OpenStudy (anonymous):

Solve the following equation for x. Identify any extraneous solutions. Show all work. 1. square root sign 8x+1=5 Solve the following equation for x. Identify any extraneous solutions. Show all work. 1. square root sign 8x+1=5 @Mathematics

OpenStudy (anonymous):

i just need to know what it means by identify any extraneous solutions.

OpenStudy (anonymous):

i don't know

OpenStudy (anonymous):

do you have selected answer?

OpenStudy (anonymous):

no. i have a bunch of questions but i have to find the answer myself.

OpenStudy (anonymous):

is this like this? 1. square root(8x+1)=5

OpenStudy (anonymous):

yeah but no parenthesis.

OpenStudy (anonymous):

1. square root sign (8x+1)=5 squaring both sides gives 8x+1=5^2=25 8x=25-1 8x=24 x=24/8=3 ans

OpenStudy (anonymous):

no parenthesis? is the square root only in 8x?

OpenStudy (anonymous):

it's sqrt sign 8x+1=5

OpenStudy (anonymous):

it's solving radical equations.

OpenStudy (anonymous):

1. square root(8x)+1=5 square root(8x)=5-1 square root(8x)=4, squaring both sides gives 8x=4^2 8x=16 x=16/8=2 ans

OpenStudy (anonymous):

any more question?